# Unifying Demonstration Selection and Compression for In-Context Learning

### v1

@@ Abstract
In-context learning (ICL) facilitates large language models (LLMs) exhibiting spectacular emergent capabilities in various scenarios. Unfortunately, introducing demonstrations easily makes the prompt length explode, bringing a significant burden to hardware. In addition, random demonstrations usually achieve limited improvements in ICL, necessitating demonstration selection among accessible candidates. Previous studies introduce extra modules to perform demonstration compression or selection independently. In this paper, we propose an ICL framework UniICL, which Unifies demonstration selection and compression, and final response generation via a single frozen LLM. Specifically, UniICL first projects actual demonstrations and inference text inputs into short virtual tokens, respectively. Then, virtual tokens are applied to select suitable demonstrations by measuring semantic similarity within latent space among candidate demonstrations and inference input. Finally, inference text inputs together with selected virtual demonstrations are fed into the same frozen LLM for response generation. Notably, UniICL is a parameter-efficient framework that only contains 17M trainable parameters originating from the projection layer. We conduct experiments and analysis over in- and out-domain datasets of both generative and understanding tasks, encompassing ICL scenarios with plentiful and limited demonstration candidates. Results show that UniICL effectively unifies 12×12$\times$12 × compression, demonstration selection, and response generation, efficiently scaling up the baseline from 4-shot to 64-shot ICL in IMDb with 24 GB CUDA allocation [^1] The code and model will be released in the final version..
@@
Jun Gao School of Computer Science and Technology Soochow University junegao1106@gmail.com
Figure 1: The efficiency comparison between UniICL and other compression methods in IMDb with the number of shots increasing from 0 to 64. * and the break of line chart indicates occurring memory explodes.
Figure 2: Compressing candidate demonstrations into spans of compressed virtual tokens and selecting one for the inference inputs. Compressed virtual tokens substitute the original demonstration together with the inference input, fed to the target LLM for prediction generation.
@@ 1 Introduction
In-context learning (ICL) Brown et al. (2020); Xie et al. (2021); Wang et al. (2023b) exhibits powerful performance in practical applications with the emergence of scaled-up Transformer-based Vaswani et al. (2017) Large Language Models (LLMs) Wang et al. (2023c); Yang et al. (2023); Wei et al. (2023); Wang et al. (2023a); Min et al. (2022). In ICL, the provided demonstrations activate the pre-training knowledge of LLMs to allow them to perform well on various downstream tasks such as text summarization Wang et al. (2023c); Yang et al. (2023) and text classification Min et al. (2022) without gradient updating of billion parameters. Despite its significant role in the era of LLMs, ICL also brings an enormous challenge to the input window. Specifically, inevitably introducing demonstrations directly causes length disaster Wang et al. (2024), which is more serious in long generative tasks such as question answering Ghosal et al. (2022) and text summarization Narayan et al. (2018), bringing significant memory costs and decreasing inference throughput as described in Figure 1. Except for length disaster, Liu et al. (2021) points out that the quality of selected demonstrations significantly influences the ICL performance. The mentioned two issues derive two research lines focused on selecting salient demonstrations and utilizing demonstrations efficiently.
Generally, researchers attempted to alleviate length explode Wingate et al. (2022); Mu et al. (2023); Ge et al. (2023) via soft prompts. The main idea was to train an independent compressor to compress the input into compact virtual tokens. Similarly, Liu et al. (2021); Lu et al. (2021) scores each candidate demonstration for the given inference input via an extra ranker, and Ram et al. (2023); Wang et al. (2024) finds that a fine-tuned LLM is up to the task of ranking. However, the external compressor/ranker naturally leads to the Out-of-Distribution (OoD) problem between the compressor/ranker and the target LLM Jiang et al. (2023), requiring extra large-scale data to perform alignment. In addition, the standalone compressor or ranker incurs additional memory costs as it necessitates simultaneous loading alongside the target LLM.
Motivated by the inherent semantic understanding ability of LLMs, we propose an ICL framework UniICL that unifies demonstration compression, demonstration selection, and response generation via the same frozen LLM. In UniICL, the same frozen LLM is applied thrice: it acts as the compressor to compress demonstrations into virtual tokens, as the selector to select suitable demonstrations based on the saliency measured by virtual tokens, and as the generator to produce final responses under the guidance of selected virtual demonstrations. Specifically, similar to the mask token in pretrained language understanding models Devlin et al. (2018); Liu et al. (2019), UniICL introduces the compression slot [M] and attaches $k\times[\text{M}]$ behind demonstrations to be condensed. The LLM forwards each processed demonstration and outputs corresponding $k$ hidden states on top of compression slots. Considering the latter compression slots naturally summarize the preceding demonstrations due to the transformer blocks, we train a simple projection layer to convert continuations to condensed virtual tokens. We briefly explain UniICL on the linguistic acceptability task concerning how it selects one demonstration among two compressed candidates and produces the prediction in Figure 2. On the one hand, virtual tokens effectively absorb semantic information from actual demonstrations, along with inference inputs, fed into the target LLM to perform generation as normal ICL. On the other hand, virtual tokens are again applied to measure saliency scores between demonstrations and given inference inputs within latent space, while demonstrations with higher scores are preferentially fed to the generator for response generation. Furthermore, UniICL compresses each demonstration independently of others, allowing caching virtual token sequences to configure a Demonstrations Bank (DB) for reusing, which will significantly bolster the inference throughput as demonstrated in Figure 1.
Generally, since the compressor, selector, and generator are the same frozen LLM, UniICL merely equips 17M trainable parameters originating from a projection layer and one special embedding [M], and we collect 30k public data for training. We extensively evaluate UniICL on in- and out-domain datasets of both generative and understanding tasks, covering the mainstream high- and low-resource ICL [^2] High- and low-resource ICL are distinguished by the magnitude of candidates. The former usually selects demonstrations from plentiful candidates, and the latter merely selects from a handful of data.. Results indicate that UniICL effectively substitutes 12$\times$ longer demonstrations with virtual tokens and scale up the inner LLM from 4-shot to 64-shot, significantly alleviating length disaster at an expense of 8% extra inference latency (without caching). Meanwhile, compared with the popular dense retriever, UniICL selects more suitable demonstrations in different tasks.
Our main contributions are as follows:
@@
  • To our knowledge, we are the first to propose an ICL framework that unifies compression, selection, and generation via a single frozen LLM, inherently bypassing the OoD problem.
@@
  • Simple application of virtual tokens outperforms the popular dense retriever in selecting suitable demonstrations after jointly optimized by compression and selection augmentation loss.
@@ 2 Related Work
In this section, we simply rethink the methods focus on demonstration (prompt) compression and document retrieval.
@@ 2.1 Soft Prompt Compression
Recently, researchers attempted to utilize soft prompts to convert actual tokens to dense-information virtual tokens. Mostly from a distillation perspective, Wingate et al. (2022) aligned the teacher model and the student model, where the teacher model accepted the actual task instruction while the student model fed the soft prompt. The main drawback of this approach was the lack of generalization that necessitated training for each lexically different instruction. To tackle the generalization problem, Mu et al. (2023) proposed to learn a Llama-7b to compress instruction to virtual tokens, but only compress instruction was not powerful enough since the demonstrations were much longer in practical. To compress the demonstrations, Chevalier et al. (2023) proposed AutoCompressor to generate compressed virtual tokens based on a fine-tuned LLM Zhang et al. (2022). They first randomized segmented the texts with thousands of words into model-accepted range and then recursively generated soft prompts for each segment. Similarly, Ge et al. (2023) proposed ICAE that employed a LoRA-adopted Llama-7b Touvron et al. (2023) to compress the processed demonstrations to compact virtual tokens, and then fed to a frozen LLM to perform generation. Unlike previous works, the underlying LLM in UniICL was kept frozen during training and was trained on only 30k data.
@@ 2.2 Extractive Compression
3 Methodology
We propose UniICL, a parameter-efficient ICL framework that unifies demonstration compression, demonstration selection, and response generation via a single LLM. Specifically, UniICL applies a general frozen LLM thrice, as a compressor to compress demonstrations into short virtual tokens, as a selector to choose suitable demonstrations for ICL, and as a generator to produce responses. As for the selection of the underlying LLM, previous work has proposed that the Decoder-only model performs better than the Encoder-Decoder model in prompt compression Mu et al. (2023). We follow this conclusion and employ popular Vicuna-7b Zheng et al. (2023) and BlueLM-7B [^3] We mainly utilize Vicuna-7B to verify the effectiveness, the experiments of BlueLM will be exhibited in Appendiex B for concise. Team (2023) as the de facto backbone in UniICL.
We present a comparison between UniICL and other recent compression methods in Table 1. Without regard to the Lora-adopted compressor, ICAE Ge et al. (2023) is most similar to UniICL in architecture and compression mode that parameter-efficiently compresses actual demonstrations into dense soft prompts.
Figure 3: Demonstration compression.
@@ 3 Methodology
3.1 Preliminary
Empirically, a prompt may encompass in-context demonstrations and an inference input in the ICL scenario. To avoid misleading and follow the existing definitions, we assume a prompt $P=(x^{D}_{1},x^{D}_{2},...,x^{D}_{n},x^{I})$, where $x^{D}_{i}$ is the $i$th demonstrations consist of an input-output pair (e.g. [DOC] -> [Sum]), and $x^{I}$ is the inference input (e.g. documents in the summarization task). We first initialize the compression slots [M] $\in\text{R}^{d}$ from the rarely used embedding of corresponding LLM, with gradient updating during training, $d$ is the dimension of the input embeddings. The compression ratio $r$ is a hyper-parameter, which is dynamically sampled from 2 to 16 during training, to calculate the specific number of compressed slots by $k=|x^{D}|/{r}$.
@@ 3.2 Demonstration Compression
For each demonstration in $P$, UniICL activates the compression slots to aggregate information from demonstrations in the forward propagation of frozen Vicuna respectively, as illustrated in Figure 3. Practically, we first attach $k$ compression slots $=k\times\text{[M]}$M] to each demonstration $x^{D}_{i}$, formatting modified prompt fed to the Vicuna. Then, frozen Vicuna forwards the modified prompts and outputs the last hidden states $H_{i}=(h_{1},h_{2},...,h_{k})$ corresponding to $k$ compression slot, dropping the others444$\oplus$ means token-level concatenation.:
_$\_,H_{i}=\text{Encode}(x^{D}_{i}\oplus M)$ (1)
Due to the transformer block in Vicuna, $H_{i}$ is compelled to attend to the preceding actual tokens and UniICL inserts a linear layer to project $H_{i}$ into compressed virtual tokens $C^{i}=(c_{1},c_{2},...,c_{k})$:
$c_{j}=W_{p}\cdot h_{j},$ (2)
@@ 3.3 Demonstration Selection
Except for substituting the origin demonstrations for generation, activated virtual tokens $C^{i}$ are again applied for demonstration selection among candidates, as illustrated in Figure 4. Specifically, given an inference input $x^{I}$ and its candidate demonstrations $(x^{D}_{1},x^{D}_{2},...,x^{D}_{n})$, UniICL obtains their latent representation after average pooling:
$\bar{C^{i}}=\frac{1}{k}\sum_{j=1}^{k}c_{j},$ (3)
denote as $q$ for the input and $(k_{1},k_{2},...,k_{n})$ for candidate demonstrations for simlifying. We define the cosine similarity between $q$ and $k_{i}$ is the $i$th demonstration saliency score $S_{i}$:
$S_{i}=cos(q,k_{i}).$ (4)
@@ 3.4 In-context Generation
We employ the frozen Vicuna again to generate responses with the guiding of concatenated virtual demonstrations, as illustrated in Figure 5. For $m$-shot in-context learning, we obtain $m$ virtual demonstrations after previously mentioned slot activating and demonstration selecting, denoted as $C^{1}$ to $C^{m}$. Then, we horizontally concatenate them, keeping their relative position unmodified. Finally, the concatenated virtual demonstrations together with actual inference inputs are fed into Vicuna, performing auto-regressive generation as normal:
$y_{i}=\text{Decode}(C^{1},...,C^{m};x^{I};y_{<i})$ (5)
Except for the generative manner, virtual tokens are conveniently transferred to close-ended evaluation without modifications for understanding tasks through providing the candidate answers and measuring their perplexity (PPL)  [^5] https://huggingface.co/docs/transformers/perplexity respectively, e.g. $(ppl^{+},ppl^{-}$) for sentiment classification:
$\text{Prediction}=argmin(ppl^{+},ppl^{-}),$ (6)
@@ 3.5 Training
The trainable parameters in UniICL are merely 17M originating from the projection layer $W_{p}$ and the introduced compression slot [M]. We first use the common language modeling loss to optimize UniICL and get a base compression model. Then we add InfoNCE He et al. (2020) to augment the demonstration selection ability:
@@
Specifically, during the first training phase, we slice the source input of each training instance into two parts and randomly compress one, denoted the compression part as $x_{c}$ and the unmodified as $x_{u}$, respectively. Afterward, we attach $M$ to $x_{c}$ and get virtual tokens $C$ on top of the compression slot sequence, as described in Equ. 1 and Equ. 2. Therefore, the language modeling loss $\mathcal{L}_{lm}$ if obtained as:
$\mathcal{L}_{lm}=-\frac{1}{|y|}\sum_{t=0}logP(y_{t}|x_{2};C;y_{<t}),$ (8)
where $y=(y_{1},y_{2},...)$ is the reference label of the current training instasnce. Notably, distinguished from ICAE which always puts soft prompts in the front, UniICL keeps the relative position of $C$ and $x_{u}$ unchanged to make compressed tokens insensitive to their position in prompts. Namely, compressed tokens will be placed in the former (latter) if $x_{c}$ is in front (behind) of $x_{u}$ in the original sequence. Additionally, aiming to concatenate different virtual token sequences from different demonstrations for practical employment of ICL, we introduce concatenation compression that evenly slices the extra-long compression parts into two sub-segments during training, compresses them separately, and concatenates their virtual tokens horizontally.
Consequently, we utilize contrastive learning for selection augmentation and mine positives $x^{D}_{+}$ and negatives $x^{D}_{-}$ as illustrated in Figure 6. Specifically, given each training instance $x^{I}$ and n candidate demonstrations $(x^{D}_{1},x^{D}_{2},...,x^{D}_{n})$ from two non-crossing training subsets, we employ Vicuna to calculate the PPL concerning the golden label of $x^{I}$, denoted as $ppl^{I}$. Then, we provide the $i$th demonstration and calculate its PPL, denoted as $(ppl^{D}_{i},i\in[1,n])$. We count $ppl^{I}$ as the baseline and calculate candidate relative PPL gains:
$\widetilde{ppl^{D}_{i}}=ppl^{I}-ppl^{D}_{i},i\in[1,n].$ (9)
Hence, $k^{+}$ $(k^{-})$ is the latent representation of demonstrations that furthest reduces (increases) $ppl^{I}$ as processed in Equation 3, and the contrastive loss $\mathcal{L}_{ctr}$ can be formulated as:
$\mathcal{L}_{ctr}=\frac{exp(cos(q,k^{+}))}{exp(cos(q,k^{+}))+exp(cos(q,k^{-}))}.$ (10)
@@
Notably, the inner compression ratio $r$ for each training step is dynamic and randomly sampled from [2, 16] during training. After the two training phases, we blend 2,000 instances from the in-domain validation sets to evaluate the quality of virtual demonstrations provided to LLMs under various ratios. Ultimately, the preferred scheme involves compressing the demonstrations with $r=12$, more details refer to Figure 7.

### v2

@@ Abstract
In-context learning (ICL) enables LLMs to exhibit spectacular emergent capabilities in various scenarios. Unfortunately, introducing demonstrations easily makes the prompt length explode, bringing a significant burden to hardware. In addition, random demonstrations usually achieve limited improvements in ICL, necessitating demonstration selection among accessible candidates. Previous studies introduce extra modules to perform demonstration compression or selection independently. In this paper, we propose an ICL framework UniICL, which Unifies demonstration selection and compression, and final response generation via a single frozen LLM. UniICL leverages the understanding ability of well-trained LLMs to independently compress different demonstrations into compressed features, and then a learnable projection layer converts features to LLM-acceptable compressed virtual tokens. Apart from substituting original demonstrations to reduce input length, virtual tokens are again used to select potential demonstrations. Finally, current queries together with selected compressed virtual tokens are fed into the same frozen LLM for response generation. UniICL is a parameter-efficient framework that only contains 17M trainable parameters originating from the projection layer and a learnable embedding. We build UniICL upon two backbones and conduct experiments over in- and out-domain datasets of both generative and understanding tasks, encompassing ICL scenarios with plentiful and limited demonstration candidates. Results show that UniICL effectively unifies 12×12$\times$12 × compression, demonstration selection, and response generation, efficiently scaling up the baseline from 4-shot to 64-shot ICL with 24 GB CUDA allocation [^1] The code and model will be released in the final version..
@@
Jun Gao1, Ziqiang Cao1∗, Wenjie Li2 School of Computer Science and Technology1 Department of Computer Science, Hong Kong Polytechnic University2
Figure 1: UniICL compresses candidate demonstrations into spans of compressed virtual tokens and selects demonstrations similar to the query. Virtual tokens substitute the original demonstration together with the query, fed to the same LLM for prediction generation.
@@ 1 Introduction
In-context learning (ICL) Brown et al. (2020); Xie et al. (2021); Wang et al. (2023b) exhibits powerful performance in practical applications with the emergence of scaled-up Transformer-based Vaswani et al. (2017) Large Language Models (LLMs) Wang et al. (2023c); Yang et al. (2023); Wei et al. (2023); Wang et al. (2023a); Min et al. (2022). In ICL, the provided demonstrations activate the pre-training knowledge of LLMs to allow them to perform well on various downstream tasks such as text summarization Wang et al. (2023c); Yang et al. (2023) and text classification Min et al. (2022) without gradient updating of billion parameters. Despite its significant role in the era of LLMs, ICL also brings an enormous challenge to the input window. Specifically, inevitably introducing demonstrations directly causes length explosion Wang et al. (2024), bringing significant memory costs and decreasing inference throughput as described in Figure 8. Except for length explosion, Liu et al. (2021) points out that the quality of the selected demonstrations significantly influences the ICL performance.
Recent efforts in modifying model architecture significantly expand the input window Zheng et al. (2022); Wu et al. (2022); Ding et al. (2023); Bulatov et al. (2023). However, LLMs with million context windows still struggle to overcome performance degradation Liu et al. (2024). Researchers also attempted to alleviate length explosion via prompt pruning Jiang et al. (2023) or soft prompts Wingate et al. (2022); Mu et al. (2023); Ge et al. (2023). Their main idea is to train an independent compressor to compress the input into soft prompts. While in selecting powerful in-context demonstrations, Liu et al. (2021); Lu et al. (2021) scores each candidate demonstration for the current queries via an extra ranker, and Ram et al. (2023); Wang et al. (2024) finds that a fine-tuned LLM is up to the task of ranking.
However, the standalone compressor or ranker inevitably incurs additional memory costs as it necessitates simultaneous loading alongside the target LLM. To tackle these challenges, we propose a framework UniICL as illustrated in Figure 1, which leverages the semantic understanding ability of target LLMs developed during pre-training to compress and select demonstrations. UniICL keeps the target LLM frozen during training to avoid catastrophic forgetting and reduce training costs. Considering the frozen LLM within UniICL function as both compressor and generator, UniICL is training efficiently and does not need to load extra compression modules, significantly reducing memory costs for both training and inference. As the underlying LLMs in visual language models fail to understand visual features without an adapter, UniICL still requires converting compressed features into LLM-acceptable compressed virtual tokens. Apart from substituting lengthy demonstrations, virtual tokens are dense representations of original demonstrations, which can be naturally further applied to select potential demonstrations by measuring similarity. In this paper, the connector is a projection layer supervised-tuned under the language modeling objective of the generator during learning compression and jointly optimized by language modeling and contrastive objectives He et al. (2020) during learning selection.
UniICL notices the fact that demonstrations in the ICL are independent of each other, which is ignored by previous compression studies Mu et al. (2023); Wingate et al. (2022); Ge et al. (2023). Therefore, UniICL independently compresses demonstrations and proposes to configure Demonstration Bank (DB) to cache compressed virtual tokens for further reusing without requiring repeated compression for the same demonstration. We evaluate UniICL which only contains 17M trainable parameters on a scope of benchmarks involving linguistic acceptability, semantic classification, text summarization, and passage reranking, and UniICL achieves outstanding performances. Then, the in-domain experiments show that UniICL built upon two backbones (Vicuna and BlueLM) effectively substitutes the 12$\times$ demonstrations with soft prompts, easily scaling up the inner LLM from 4-shot to 64-shot. Our main contributions are as follows:
@@
  • To our knowledge, we are the first to propose an ICL framework that unifies compression, selection, and generation via a single frozen LLM.
@@
  • UniICL proposes to configure Demonstration Bank to enhance ICL efficiency, avoiding repeated compression for the same demonstration.
@@ 2.1 Soft Prompt Compression
Recently, researchers attempted to utilize soft prompts to convert actual tokens to dense-information virtual tokens. Mostly from a distillation perspective, Wingate et al. (2022) aligned the teacher model and the student model, where the teacher model accepted the actual task instruction while the student model fed the soft prompt. The main drawback of this approach was the lack of generalization that necessitated training for each lexically different instruction. To tackle the generalization problem, Mu et al. (2023) proposed to learn a Llama-7b to compress instruction to virtual tokens, but only compress instruction was not powerful enough since the demonstrations were much longer in practice. To compress the demonstrations, Chevalier et al. (2023) proposed AutoCompressor to recurrently generate compressed virtual tokens based on a fine-tuned Llama Zhang et al. (2022). However, AutoCompressor broke the independence of demonstrations, and the recurrent compression increased inference latency. Ge et al. (2023) proposed ICAE that employed a LoRA-adopted Llama-7b Touvron et al. (2023) to compress the processed demonstrations to compact virtual tokens, while ICAE still struggled to overcome quite long inputs.
@@ 2.2 Extractive Compression
3 Methodology
We propose UniICL, a parameter-efficient ICL framework that unifies demonstration compression, demonstration selection, and response generation via a single LLM. As for the selection of the underlying LLM, previous work has proved that the Decoder-only model performs better than the Encoder-Decoder model in prompt compression Mu et al. (2023). We follow this conclusion and employ popular Vicuna-7b Zheng et al. (2023) and BlueLM-7B [^2] We mainly discuss UniICL built on Vicuna-7B, the experiments of BlueLM will be exhibited in Appendiex B. Team (2023) as the underlying backbone in UniICL.
Figure 2: Differences of compression methods in formulating compressed virtual tokens in ICL.
We present a comparison of training costs between UniICL and other recent compression methods in Table 1. Additionally, we illustrate differences in formulating virtual tokens for compression methods based on the soft prompt in Figure 2.
To explain plainly, we ideally assume the compressor within three compression methods based on soft prompts has the window limitation of $L$, and has the same compression ratio, ignoring the length of soft prompts. In the 2-shot scenario, demonstrations $D_{1}$ and $D_{2}$ have a length of $L$. Considering the AutoCompressor, the concatenated demonstrations will be divided back into two segments, and the AutoCompressor compresses each segment step-by-step, bringing two times non-parallel compression. When it comes to ICAE, merely $D_{1}$ is accessible for the compressor and others will be read by no means. AutoCompressor shows advantages in the readable prompt length, but is short in efficiency, while ICAE has a constant compression complexity but struggles to approach relatively long inputs. Additionally, AutoCompressor makes the compression of later demonstrations be conditioned on the previous ones, which breaks demonstration independence in ICL.
Combining the advantages of AutoCompressor and ICAE, UniICL compresses demonstrations independently and introduces virtual tokens concatenation to overcome long prompt challenges. In the $N$-shot settings, the number of practical compression steps can be calculated as $\lceil\frac{N}{k}\rceil$, where $k$ indicates that a single GPU is capable of compressing $k$ demonstrations in a batch. When the GPU capacity is sufficient, k equals $N$, which is the scenario of ICAE that compresses all segments in a time but UniICL drops nothing, while it degenerates to the AutoCompressor scenario that compresses segments step-by-step, when the GPU capacity is only sufficient to set $k=1$.
Figure 3: Demonstration compression.
@@ 3.1 Demonstration Compression
UniICL introduces memory slots [M], a learnable embedding initialized from a rarely used embedding of the target LLM. UniICL activates the compression slots to absorb information from demonstrations in the forward propagation of frozen Vicuna, as illustrated in Figure 3. We first attach $k$ compression slots $=k\times\text{[M]}$M] to each demonstration $D_{i}$, formatting modified prompt fed to the Vicuna. Then, frozen Vicuna forwards the modified prompts and outputs the last hidden states $H_{i}=(h_{1},h_{2},...,h_{k})$ on top of the $k$ compression slots, dropping the others333$\oplus$ means token-level concatenation.:
_$\_,H^{i}=\mathrm{forward}(D_{i}\oplus M)$ (1)
Due to the attention mechanism, $H^{i}$ is compelled to attend to the preceding actual tokens. Then, UniICL inserts a projection layer to convert $H^{i}$ into LLM-acceptable compressed virtual tokens $C^{i}=(c^{i}_{1},c^{i}_{2},...,c^{i}_{k})$:
$c^{i}_{j}=W_{p}\cdot h^{i}_{j},$ (2)
@@ 3.2 Demonstration Selection
Except for substituting the origin demonstrations for generation, compressed virtual tokens $C^{i}$ are representative of demonstrations that can be again applied for demonstration selection, as illustrated in Figure 4. Specifically, given a query $Q$ and its candidate demonstrations $(D_{1},D_{2},...,D_{n})$, UniICL obtains their latent representation after average pooling:
$\bar{C^{i}}_{Q/D}=\frac{1}{k}\sum_{j=1}^{k}c_{j}.$ (3)
We define the $i$-th demonstration saliency score $S_{i}$ via the cosine similarity between Query and demonstrations:
$S_{i}=\mathrm{cos}(\bar{C_{Q}},\bar{C}_{D}^{i}).$ (4)
Demonstrations are then reranked according to their saliency scores.
@@ 3.3 In-context Generation
We employ the frozen Vicuna again to generate responses with the guiding of concatenated virtual demonstrations and queries, as illustrated in Figure 5. For $m$-shot in-context learning, we obtain $m$ virtual tokens after demonstration compression and selection, denoted as $C^{1}$ to $C^{m}$. Then, we horizontally concatenate them, keeping their relative position unmodified. Finally, the concatenated virtual tokens together with actual inference inputs are fed into Vicuna, performing auto-regressive generation as normal:
$y_{i}=\mathrm{generate}(C^{1},...,C^{m};Q;y_{<i})$ (5)
Except for the generative manner, virtual tokens are conveniently transferred to close-ended evaluation for understanding tasks through providing the candidate answers and measuring the label space PPL  [^4] https://huggingface.co/docs/transformers/perplexity, e.g. $(ppl^{+},ppl^{-}$) for sentiment classification:
$y=\mathrm{argmin}(ppl^{+},ppl^{-}),$ (6)
@@ 3.4 Training
The trainable parameters in UniICL are merely 17M originating from the projection layer $W_{p}$ and the introduced compression slot [M]. The projection layer is optimized with the language modeling objective of Vicuna to learn a base compression model. Then InfoNCE He et al. (2020) joint with language modeling objective are used to augment the demonstration selection ability of the base compression model:
@@
Specifically, we slice the source input of each training instance into two parts and randomly compress one, denoted the compression part as $x_{c}$ and the unmodified as $x_{u}$. Afterward, we attach $M$ to $x_{c}$ and get virtual tokens $C$ on top of the compression slots, as described in Equ. 1 and Equ. 2. Therefore, the language modeling loss $\mathcal{L}_{lm}$ is obtained as:
$\mathcal{L}_{lm}=-\frac{1}{|y|}\sum_{t=0}logP(y_{t}|Q;C;y_{<t}),$ (8)
where $y$ is the reference label of the current training instance. Additionally, to approach the large-shot settings without significant truncation, we introduce concatenation compression. When $x_{c}$ exceeds the window limitation for compression, UniICL further divides $x_{c}$ into acceptable ranges and compresses them independently to get local virtual tokens. Then, these virtual tokens of several segments will be concatenated to formulate global virtual tokens to replace $x_{c}$.
Consequently, we utilize contrastive learning for selection augmentation and mine positives and negatives as illustrated in Figure 6. Specifically, given each training instance $Q$ and $n$ candidate demonstrations $(D_{1},D_{2},...,D_{n})$ from two non-crossing training subsets, we employ Vicuna to calculate the PPL concerning the golden label of $Q$, denoted as $ppl^{Q}$. Then, we provide the $i$-th demonstration and calculate PPL concerning the golden label of $Q$, denoted as $(ppl^{D}_{i},i\in[1,n])$. We count $ppl^{Q}$ as the baseline and calculate candidate relative PPL gains:
$\widetilde{ppl^{D}_{i}}=ppl^{Q}-ppl^{D}_{i},i\in[1,n].$ (9)
Hence, $C_{D}^{+}$ $(C_{D}^{-})$ is the representation of demonstrations $D^{+}$ ($D^{-}$) that furthest reduces (increases) $ppl^{Q}$ as processed in Equ. 3, and the contrastive loss $\mathcal{L}_{ctr}$ can be formulated as:
$\mathcal{L}_{ctr}=\frac{exp(cos(C_{Q},C_{D}^{+}))}{exp(cos(C_{Q},C_{D}^{+}))+exp(cos(C_{Q},C_{D}^{-}))}.$ (10)